OpenStudy (anonymous):

I have done A. and gotten 1 MB = 2^23 Bits I need some help on B. Please explain to me what to do and how to get the answer for B. A. Write each answer as a power of 2. Show your work and explain your steps. a. Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or 0, in the computer's memory. A byte is 2^3 bits. A megabyte (MB) is 2^20 bytes. How many bits are in a megabyte? B. A gigabyte (GB) is 2^30 bytes. How many Megabytes are in a gigabyte? How many bits are in a gigabyte?

OpenStudy (anonymous):

Okay, I could be wrong on this, but I'll give it my thoughts. If a GB is 2^30 Bytes, and a MB is 2^20 Bytes, then that means there are 1048576 Bytes in a MB and 1073741824 Bytes in a GB. (2^20)/(2^30)=9.765625E-4. That would make a MB 0.000976562 of a GB. 1/0.000976562=1024 , so there are 1024 MB in a GB. As for the second part, it's just (2^3)*(2^30).

OpenStudy (debbieg):

Don't let conversions scare you. Conversions are just a matter of lining up the necessary conversion factors to switch the units. Each conversion factor is just a fraction where the num'r = the den'r. Just make sure the units are where you need them, e.g., whatever units you start with will cancel with something in the den'r of the converson factor and leave you with the units in the num'r. So for example, 12 in = 1 ft. Thus I can use a conversion factor of \(\Large \dfrac{ 12 ~\text{in} }{ 1~\text{ ft} }\) or \(\Large \dfrac{ 1 ~\text{ft} }{ 12~\text{ in} }\) depending on the "direction of the conversion. so how many INCHES are in 3 FT? I'm starting with ft, I want to end up with inches, so: \(\Large 3~\text{ft}\cdot\dfrac{ 12 ~\text{in} }{ 1~\text{ ft} }\) \(=\Large 3~\cancel {\text{ft}}\cdot\dfrac{ 12 ~\text{in} }{ 1~\cancel {\text{ ft} }}=36 in\)

OpenStudy (debbieg):

Now apply that to your problem with bits and bytes and MB and GB, etc..... so you'll have conversion factors like:

OpenStudy (anonymous):

Someone else helped me with this earlier guys but thanks for the help

OpenStudy (debbieg):

\(\Large \dfrac{ 1 ~\text{MB} }{ 2^{20}~\text{bytes} }\) \(\Large \dfrac{ 1 ~\text{GB} }{ 2^{30}~\text{bytes} }\)